OpenStudy (smartbrain):

CDEF is a rectangle with a width of 3 and a length of 4. One diagonal is 3x - 2y and the other diagonal is 8x + 3y. Find the values of x and y.

OpenStudy (smartbrain):

@ParthKohli @pooja195 @Michele_Laino Can one of you guys help me? Am I doing this right?

OpenStudy (smartbrain):

8x + 3y = 3x - 2y - 2y =1 8x + 1 = 3x - 3x = 5x y = 1 x = 5 ???

OpenStudy (smartbrain):

The answer choices are... x = 1 and y = 2 x = 1 and y = -1 x = 2 and y = 1 x = 1 and y = -2 So I know I am wrong.

OpenStudy (anonymous):

You started right. The diagonals are equal, but you need 2 equations since you have two variables. You can use the Pythagorean theorem/triples to find the length of the diagonal as a number

OpenStudy (smartbrain):

When I did that I got 3, 4, 5 (so 5 will be the missing triples)

OpenStudy (anonymous):

right, so now you can write a system of equations. \(3x-2y=5\) \(8x+3y=5\)

OpenStudy (smartbrain):

Now I combine them into one? 3+8 = 11x + 1y = 5?

OpenStudy (anonymous):

no. solve using elimination

OpenStudy (anonymous):

combining into 1 won't help yet because you have 2 variables. The y's have opposite signs, so then can be eliminated. Multiply the top equation by 3 and the bottom by 2.

OpenStudy (smartbrain):

3x - 2y = 5 * 3 = 3x - 6y = 5 8x + 3y = 5 *2 = 8x + 6y = 5 Is this right?

OpenStudy (smartbrain):

3x - 6y = 5 8x + 6y = 5 11x = 5 5 / 11 = x x = 0.45 ?

OpenStudy (anonymous):

not quite. Multiply the WHOLE equation (both sides) The first equation multiplied by 3: \[3(3x -2y) = 3(5)\] becomes \[9x-6y=15\]

OpenStudy (anonymous):

What's the 2nd multiplied by 2?

OpenStudy (smartbrain):

8x + 3y = 5 *2 = 16x + 6y = 10

OpenStudy (anonymous):

Yes. Now add the two equations